Solve 3+2+3×2/√3× 2/√3 × 2/√3=9
1
Expert's answer
2019-04-05T09:43:17-0400

Solution

3+2+3×2/√3× 2/√3 = 5+12/3=5+4=9


Answer

3+2+3×2/√3× 2/√3 =9